OpenStudy (anonymous):

factor completly step by step 3x^3+24

OpenStudy (apoorvk):

'3' is the only common factor to both the terms in the expression above.so take it out of the brackets. what do you think it ll be??

OpenStudy (anonymous):

Use the following identity:\[a^3 + b^3 = (a + b) (a^2 - a b + b^2)\]\[3\left(x^3+8\right)=3 (x+2) \left(x^2-2 x+4\right) \]